OpenStudy (anonymous):

how will we know the wavefunction which is normalized for time t=0 is also normalized for all times?

OpenStudy (abb0t):

There is an integral function that defines this. I am sure you can find this in your book: \[\frac{ d }{ dt } \int\limits |\psi (x,t)|^2dx=\frac{ i \cal{h} }{ 2m }[\psi^*\frac{ \partial\psi }{ \partial x }-\frac{ \partial \psi^* }{ \partial x }\psi ]\] Now the right hand side needs to be shown to be zero but this is only true if we make the assumption that \(\psi\) is normalizable at all time, t. However, you can find functions which change from normalizable to non normalizable and still satisfy the schrodinger Equation.

OpenStudy (abb0t):

btw, that \(h\) is the hamiltonian operator. I didn't know how to change it using LaTex.

OpenStudy (abb0t):

For example, if ur given a 1-D finite square well potential, \(V_0\). The spectrum of the Hamiltonian, \(h\) includes one or more (normalizable) bound states with \(E < 0\), and a continuum of un-normalizable scattering states with \(E > 0\). Suppose initially the system is in a bound state with energy \(E_0 < 0\). Then, at \(t=0\), you turn on an additional time-varying potential \(V_1\) which oscillates at frequency \(\omega\). And \(E_0 + h\omega > 0\), so this gives the system enough energy to escape the well to infinity. In fact since \(V_1\) includes a step function at \(t = 0\), it contains a continuous range of frequencies, and as a result the system's energy will also have a continuous range. In terms of perturbation theory, \(V_1\) is a perturbation connecting a normalizable state with an unormalizable one, and the system has evolved from one to the other.